OpenStudy (anonymous):

name properties that justify this statement. all variables represent real #'s : ( -2.6 ) + 7 = 7 + ( -2.6 ) and y . 1/y = 1

OpenStudy (anonymous):

name properties that justify this statement. all variables represent real #'s : ( -2.6 ) + 7 = 7 + ( -2.6 ) and y . 1/y = 1